# Quench Spectroscopy for Dissipative and Non-Hermitian Quantum Lattice Models

### v1

@@ 5 Quench spectroscopy of the double-quenched Bose-Hubbard chain
In Refs. [68, 7, 8], the so-called quench spectroscopy (QS) method has been introduced theoretically whose predictions have been verified numerically using tensor-network-based techniques for one-dimensional quantum systems. The latter permits to determine the low-lying excitation spectrum or quasiparticle dispersion relation of quantum lattice models by analysing their dynamical response to a sudden quench which can be either local, i.e. the perturbation is applied on a specific lattice site, or global, i.e. all the lattice sites are concerned. For instance, it has been applied to disordered quantum lattice models  [70, 71]. In what follows, we focus on the case of sudden global quenches. The main feature of this method corresponding also to the most restrictive condition is to consider a so-called weak quench meaning that the many-body ground state of the post-quench Hamiltonian has to be relatively closed to, i.e. strongly overlaps with, the initial state corresponding the ground state of the pre-quench Hamiltonian. In other words, the method holds for sudden global quenches confined in a same gapless or gapped phase and ideally in the same regime within this quantum phase. According to the QP theory used previously and detailed in Refs. [68, 57], equal-time connected correlation functions can generally be cast into a generic form. The latter reads as:
$G(R,t)=\int_{\mathcal{B}}\mathrm{d}k\leavevmode\nobreak\ \mathcal{S}_{k}\left\{e^{i(kR+2\mathcal{E}_{k}t)}+e^{i(kR-2\mathcal{E}_{k}t)}\right\},$ (37)
@@
where $\mathcal{E}_{k}$ refers to the quasiparticle dispersion relation of the post-quench Hamiltonian. The so-called quench spectral function (QSF) denoted by $S_{k}(\omega)$ is defined as the $1+1$ (space-time) Fourier transform of the equal-time connected correlation function $G(R,t)$ and reads as  [8]:
$S_{k}(\omega)=\int_{0}^{L}\mathrm{d}R\int_{0}^{T}\mathrm{d}t\leavevmode\nobreak\ G(R,t)e^{-i(kR+\omega t)},$ (38)
@@
As expected, the previous form of the QSF clearly permits to determine the post-quench low-lying excitation spectrum $\mathcal{E}_{k}$. Indeed, from Eq. (39), $S_{k}(\omega)$ represents twice the post-quench quasiparticle dispersion relation $2\mathcal{E}_{k}$. More precisely, the latter will display the two branches $2E_{\mathbf{k}}$ and $-2E_{\mathbf{k}}$. The prefactor $2$ comes from the weak sudden global quench dynamics of the quantum lattice model being mediated by the spreading of quasiparticle pairs. The weight associated to each momentum-dependent energy is characterised by the amplitude function $\mathcal{S}_{k}$ depending on the quench parameters as well as the observable defining the equal-time correlation function $G(R,t)$. According to Eq. (39), the QSF has $k$/$-k$ symmetry; the latter coming from the post-quench quasiparticle dispersion relation $\mathcal{E}_{k}$ and is $\omega$/$-\omega$ symmetric due to the parity of the Dirac delta function $\delta$.
@@
where $G_{2}^{U}(R,t)$ and $G_{2}^{\gamma}(R,t)$ refers to the individual contribution for $G_{2}$ due to the weak sudden global quench on the repulsive interaction strength $U$ and the dissipation rate $\gamma$ respectively. The first equality is valid according to the previous findings (see Fig. 3 and the associated discussion) where both quantum quenches seem to be independent. The second one comes from the fact that the quench on $\gamma$ leads to a spatial-independent correlation spreading as depicted on Fig. 3. Indeed, from Ref. [59] where an unique quench on the dissipation strength $\gamma$ is considered, $G_{2}^{\gamma}(R,t)$ reads as (see also Eq. (35)):
$G_{2}^{\gamma}(R,t)=g_{2}^{\gamma}(R,t)-g_{2}^{\gamma}(R,0);$ (41)
$G_{2}^{\gamma}(R,t)=n(t)^{2}+\frac{2n(t)}{L}\sum_{k\neq 0}\cos(kR)\left[G_{k}(t)+\operatorname{Re}(F_{k}(t))\right]-n(0)^{2}-\frac{2n(0)}{L}\sum_{k\neq 0}\cos(kR)\left[G_{k}(0)+\operatorname{Re}(F_{k}(0))\right].$ (42)
For the BH chain in the SF-mean-field, the first excitations having a momentum $k=0^{+}$ are the relevant ones. This implies that the associated wavelength $\lambda$ is much higher than the lattice spacing $a$, i.e. $\lambda\gg a$. Hence, it follows that the discretization of the lattice is irrelevant and thus that the contribution $G_{2}^{\gamma}(R,t)$ is spatially independent, i.e. $G_{2}^{\gamma}(R,t)=G_{2}^{\gamma}(0,t)$. This can be verified straightforwardly by injecting $k=0^{+}$ in Eq. (40). Finally, it immediately follows from Eq. (40) that:
@@
Therefore, the consequence of the quench on $\gamma$ for the quench spectral function is minimal according to Eq. (43). Indeed, the latter will increase the amplitude of $S_{k}(\omega)$ for the quasimomentum $k=0$ for any energy $\omega$. This effect is shown on Fig. 4. On Fig. 4, the QSFs associated to the $G_{2}$ and $G_{1}$ ETCCF are displayed and are benchmarked with twice the post-quench quasiparticle dispersion relation associated to the isolated BH chain confined in the SF-mean-field regime, i.e. $2\mathcal{E}_{k}$, having the following expression (see also Eq. (18)):
@@
where $\bar{n}=N(0)/L$ refers here to the initial filling of the lattice chain where $N(0)$ denotes the initial total number of bosonic particles on the lattice. Note that for $G_{1}$, the spatial dependence of the contribution due to the quench on the dissipations, i.e. $G_{1}^{\gamma}(R,t)$, becomes slightly relevant as suggested on Fig. 3. This results in non-zero amplitudes for the associated QSF for a momentum $k$ in the limit $k\rightarrow 0$ and for any energy $\omega$. The latter is consistent with the fact that the first excitations have a momentum close to zero.
Figure 4: Quench spectral function $S_{k}(\omega)$ associated to (left) $G_{2}(R,t)=\langle\hat{n}_{R}(t)\hat{n}_{0}(t)\rangle_{\mathrm{c}}$ the equal-time connected density-density correlation function (right) $G_{1}(R,t)=\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{t}-\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{0}$ the equal-time connected one-body correlation function. A double quench comprising not only a weak sudden global quench on the dissipation strength from $\gamma=0$ to $\gamma>0$ but also another one on the two-body repulsive interaction strength $U$ from $U=U_{\mathrm{i}}$ to $U=U_{\mathrm{f}}$ is considered such that the quantum system is confined in the SF phase. The dashed-dotted magenta lines represent twice and minus twice the post-quench quasiparticle dispersion relation associated to the isolated BH chain confined in the SF-mean-field regime, i.e. $2\mathcal{E}_{k,\mathrm{f}}$ and $-2\mathcal{E}_{k,\mathrm{f}}$ respectively, see Eq. (44). Note that half of the two low-lying excitation spectra are missing resulting in a loss of the $k/-k$ and $\omega/-\omega$ symmetries. Indeed, the latter would be characterized by a negative group velocity and only the correlation pattern with $R,t\geq 0$ is considered here. To recover the second half, the space-time region of the correlation pattern with $R\leq 0$ and $t\geq 0$ has to be considered. The considered parameters to plot both QSFs are: $\bar{n}=1$ with $N(0)=L=400$ to satisfy the thermodynamic limit, $J=1$, $U_{\mathrm{i}}=2$, $U_{\mathrm{f}}=0.5$, $\gamma=0.1$.
@@
Figure 5: Quench spectral function $S_{k}(\omega)$ associated to (left) $G_{2}(R,t)=\langle\hat{n}_{R}(t)\hat{n}_{0}(t)\rangle_{\mathrm{c}}$ the equal-time connected density-density correlation function (right) $G_{1}(R,t)=\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{t}-\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{0}$ the equal-time connected one-body correlation function. A weak sudden double quench is performed on the dissipation strength from $\gamma=0$ to $\gamma>0$ as well as on the two-body repulsive interaction strength $U$ from $U=U_{\mathrm{i}}$ to $U=U_{\mathrm{f}}$ such that the quantum system is confined in the SF phase. The dashed-dotted magenta lines represent twice and minus twice the post-quench quasiparticle dispersion relation associated to the isolated BH chain confined in the SF-mean-field regime, i.e. $2\mathcal{E}_{k,\mathrm{f}}$ and $-2\mathcal{E}_{k,\mathrm{f}}$ respectively, see Eq. (44). The parameters are: $\bar{n}=1$ with $N(0)=L=400$ to satisfy the thermodynamic limit, $J=1$, $U_{\mathrm{i}}=2$, $U_{\mathrm{f}}=0.5$, $\gamma=1$.
@@
We turn to the case of large post-quench interaction strength denoted by $U_{\mathrm{f}}$ while considering small dissipations, i.e. $\gamma\ll 1$. For large enough $U_{\mathrm{f}}$, the 1D BH model is no longer confined in the mean-field regime of the SF phase but in the strongly-correlated regime. Indeed, the mean-field condition given by $\bar{n}\gg U/J$ is no longer satisfied when the dimensionless parameter $U/J$ and the filling of the lattice $\bar{n}$ have approximately the same value. In what follows, we consider the specific case where $J=U_{\mathrm{f}}=\bar{n}=N(0)/L$. However, from Ref. [58], it has been shown that the quasiparticle dispersion relation $\mathcal{E}_{k}$ for the BH chain confined in the SF-mean-field regime defined at Eq. (44) remains accurate far beyond its validity domain. Hence, even at large post-quench interaction strength $U_{\mathrm{f}}$, we should be able to rely on the QS approach in order to get an accurate description of the low-lying excitation spectrum of the 1D BH model confined in the SF-strongly-correlated regime. On Fig. 6, the QSF associated to the density fluctuations via the ETCCF $G_{2}$ and to the phase fluctuations via $G_{1}$ are compared to the theoretical expression of $\mathcal{E}_{k,\mathrm{f}}$ initially valid in the SF-mean-field. A very good benchmark is found for both QSFs.
@@
Figure 6: Quench spectral function $S_{k}(\omega)$ associated to (left) $G_{2}(R,t)=\langle\hat{n}_{R}(t)\hat{n}_{0}(t)\rangle_{\mathrm{c}}$ the equal-time connected density-density correlation function (right) $G_{1}(R,t)=\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{t}-\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{0}$ the equal-time connected one-body correlation function. A weak sudden double quench is performed on the dissipation strength from $\gamma=0$ to $\gamma>0$ as well as on the two-body repulsive interaction strength $U$ from $U=U_{\mathrm{i}}$ to $U=U_{\mathrm{f}}$ such that the quantum system is confined in the strongly-correlated regime of the SF phase. The dashed-dotted magenta lines represent twice and minus twice the post-quench quasiparticle dispersion relation associated to the isolated BH chain confined in the SF-mean-field regime, i.e. $2\mathcal{E}_{k,\mathrm{f}}$ and $-2\mathcal{E}_{k,\mathrm{f}}$ respectively, see Eq. (44). The parameters are: $\bar{n}=1$ with $N(0)=L=400$ to satisfy the thermodynamic limit, $J=1$, $U_{\mathrm{i}}=2$, $U_{\mathrm{f}}=1$, $\gamma=0.1$.
@@
6 Experimental feasibility of the quench spectroscopy method
@@
  • The restrictive choice for the observables in order to compute the associated connected equal-time correlation function.
  • The different sources of noise including the shot and technical noise.
  • The difficulty for the imaging due to the coherence of the time-evolved quantum state when considering the gapless SF phase contrary to the gapped MI phase.
  • The resolution associated to the experimental data to describe accurately the causal region of the space-time correlations. The latter region is primordial to recover the full information regarding the low-lying excitation spectrum of the quantum lattice model. Indeed, reliable experimental data on the correlation edge, which is already not an easy task for experimentalists, is not sufficient. The latter contains only information on the quasiparticle pair propagating with the fastest group velocity.
  • From a more general viewpoint, a main limit of the QS method is to be valid and reliable for weak global quenches. This "quench weakness" implies difficulties for experimentalists. To illustrate the previous difficulty, a strong global quench confined in the MI phase is much easier to handle compared to a weak quench in the same phase. Indeed, for the imaging and more precisely for the mobility of the bosonic particles, a strong global quench is necessary. By "strong quench", we mean here that the initial or pre-quench dimensionless interaction parameter is significantly higher or lower than the final or post-quench dimensionless interaction parameter, i.e. $(U/J)_{\mathrm{i}}\gg(\ll)\leavevmode\nobreak\ (U/J)_{\mathrm{f}}$.
  • Another limitation of the method lies in the kind of loss processes which is considered. It is entirely possible to imagine loss processes leading to a non-trivial, i.e. strong, spatial dependence for the space-time pattern of the ETCCF.
The latter QS method to probe low-lying excitation spectra can be extended to dissipative many-body quantum fermionic and spin lattice chains, to a higher dimensionality of the lattice and to long-range interactions. Most importantly, this technique can be applied to experimental measurements obtained using quantum simulators based on ultracold atoms to engineer bosonic lattice models, neutral Rydberg atoms or trapped ions for the simulation of fermionic and spin lattice models. To conclude, QS can be seen as an alternative approach to the standard experimental techniques to measure excitation spectra such as the angle-resolved photoemission spectroscopy (ARPES)  [72] or Bragg spectroscopy [73].
7 Applicability of the quench spectroscopy to a non-Hermitian quantum lattice model

### v2

@@ II.4 Quench spectroscopy of the double-quenched Bose-Hubbard chain
In Refs. despres_these ; sanchezpalencia2019 ; villa2020 , the so-called quench spectroscopy (QS) method has been introduced theoretically whose predictions have been verified numerically using tensor-network-based techniques for one-dimensional (1D) quantum systems. The latter permits to determine the low-lying excitation spectrum or quasiparticle dispersion relation of quantum lattice models by analyzing their dynamical response to a sudden quench which can be either local, i.e. the perturbation is applied on a specific lattice site, or global, i.e. all the lattice sites are concerned. For instance, the QS has already been applied to disordered quantum lattice models thomson2021 ; thomson2021paper2 and quantum spin models menu2018 ; menu2023 ; bocini2024 . In what follows, we focus on the case of sudden global quenches. The main feature of this method corresponding also to the most restrictive condition is to consider a so-called weak quench meaning that the many-body ground state of the post-quench Hamiltonian has to be relatively closed to, i.e. strongly overlaps with, the initial state corresponding the ground state of the pre-quench Hamiltonian. In other words, the method holds for sudden global quenches confined in a same gapless or gapped phase and ideally in the same regime within this quantum phase. According to the QP theory used previously and detailed in Refs. despres_these ; despres2018 , equal-time connected correlation functions can generally be cast into a generic form. The latter reads as:
$G(R,t)=\int_{\mathcal{B}}\mathrm{d}k\mathcal{S}_{k}\left[e^{i(kR+2\mathcal{E}_{k}t)}+e^{i(kR-2\mathcal{E}_{k}t)}\right],$ (31)
@@
where $\mathcal{E}_{k}$ refers to the quasiparticle dispersion relation of the post-quench Hamiltonian. The so-called quench spectral function (QSF) denoted by $S_{k}(\omega)$ is defined as the $1+1$ (space-time) Fourier transform of the equal-time connected correlation function $G(R,t)$ and reads as villa2020 :
$S_{k}(\omega)=\int_{0}^{L}\mathrm{d}R\int_{0}^{T}\mathrm{d}tG(R,t)e^{-i(kR+\omega t)},$ (32)
@@
As expected, the previous form of the QSF clearly permits to determine the post-quench low-lying excitation spectrum $\mathcal{E}_{k}$. Indeed, from Eq. (33), $S_{k}(\omega)$ represents twice the post-quench quasiparticle dispersion relation $2\mathcal{E}_{k}$. More precisely, the latter will display the two branches $2E_{\mathbf{k}}$ and $-2E_{\mathbf{k}}$. The prefactor $2$ comes from the weak sudden global quench dynamics of the quantum lattice model being mediated by the spreading of quasiparticle pairs. The weight associated to each momentum-dependent energy is characterized by the amplitude function $\mathcal{S}_{k}$ depending on the quench parameters as well as the observable defining the equal-time correlation function $G(R,t)$. According to Eq. (33), the QSF has $k$/$-k$ symmetry; the latter coming from the post-quench quasiparticle dispersion relation $\mathcal{E}_{k}$ and is $\omega$/$-\omega$ symmetric due to the parity of the Dirac delta function $\delta$.
@@
where $G_{2}^{U}(R,t)$ and $G_{2}^{\gamma}(R,t)$ refers to the individual contribution for $G_{2}$ due to the weak sudden global quench on the repulsive interaction strength $U$ and the dissipation rate $\gamma$ respectively. The first equality is valid according to the previous findings (see Fig. 3 and the associated discussion) where both quantum quenches seem to be independent. The second one comes from the fact that the quench on $\gamma$ leads to a spatial-independent correlation spreading as depicted on Fig. 3. Indeed, from Ref. despresnew where an unique quench on the dissipation strength $\gamma$ is considered, $G_{2}^{\gamma}(R,t)$ reads as:
$G_{2}^{\gamma}(R,t)=$ $\leavevmode\nobreak\ g_{2}^{\gamma}(R,t)-g_{2}^{\gamma}(R,0);$ (35a)
$G_{2}^{\gamma}(R,t)=$ $\leavevmode\nobreak\ n(t)^{2}+\frac{2n(t)}{L}\sum_{k\neq 0}\cos(kR)\left[G_{k}(t)+\operatorname{Re}(F_{k}(t))\right]-n(0)^{2}-\frac{2n(0)}{L}\sum_{k\neq 0}\cos(kR)\left[G_{k}(0)+\operatorname{Re}(F_{k}(0))\right],$ (35b)
where $g_{2}(R,t)$ is defined at Eq. (29). For the BH chain in the SF-mean-field, the first excitations having a momentum $k=0^{+}$ are the relevant ones. This implies that the associated wavelength $\lambda$ is much higher than the lattice spacing $a$, i.e. $\lambda\gg a$. Hence, it follows that the discretization of the lattice is irrelevant and thus that the contribution $G_{2}^{\gamma}(R,t)$ is spatially independent, i.e. $G_{2}^{\gamma}(R,t)=G_{2}^{\gamma}(0,t)$. This can be verified straightforwardly by injecting $k=0^{+}$ in Eq. (34). Finally, it immediately follows from Eq. (34) that:
@@
Therefore, the consequence of the quench on $\gamma$ for the quench spectral function is minimal according to Eq. (36). Indeed, the latter will increase the amplitude of $S_{k}(\omega)$ for the quasi-momentum $k=0$ for any energy $\omega$. This effect is shown on Fig. 4. On Fig. 4, the QSFs associated to the $G_{2}$ and $G_{1}$ ETCCF are displayed and are benchmarked with twice the post-quench quasiparticle dispersion relation associated to the isolated BH chain confined in the SF-mean-field regime, i.e. $2\mathcal{E}_{k}$, having the following expression:
@@
where $\bar{n}=N(0)/L$ refers here to the initial filling of the lattice chain with $N(0)$ denoting the initial total number of bosonic particles on the lattice, see also Eq. (10). Note that for $G_{1}$, the spatial dependence of the contribution due to the quench on the dissipations, i.e. $G_{1}^{\gamma}(R,t)$, becomes slightly relevant as suggested on Fig. 3. This results in non-zero amplitudes for the associated QSF for a momentum $k$ in the limit $k\rightarrow 0$ and for any energy $\omega$. The latter is consistent with the fact that the first excitations have a momentum close to zero.
Figure 4: Quench spectral function $S_{k}(\omega)$ associated to (top panel) $G_{2}(R,t)=\langle\hat{n}_{R}(t)\hat{n}_{0}(t)\rangle_{\mathrm{c}}$ the equal-time connected density-density correlation function (bottom panel) $G_{1}(R,t)=\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{t}-\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{0}$ the equal-time connected one-body correlation function. The double quench here comprises a weak sudden global quench on the dissipation strength from $\gamma=0$ to $\gamma>0$ as well as one performed on the two-body repulsive interaction strength $U$ from $U=U_{\mathrm{i}}$ to $U=U_{\mathrm{f}}$. The dashed-dotted magenta lines represent twice and minus twice the post-quench quasiparticle dispersion relation associated to the isolated BH chain confined in the SF-mean-field regime, i.e. $2\mathcal{E}_{k,\mathrm{f}}$ and $-2\mathcal{E}_{k,\mathrm{f}}$ respectively, see Eq. (37). Note that half of the two low-lying excitation spectra are missing resulting in a loss of the $k/-k$ and $\omega/-\omega$ symmetries. Indeed, the latter would be characterized by a negative group velocity and only the correlation pattern with $R,t\geq 0$ is considered here. To recover the second half, the space-time region of the correlation pattern with $R\leq 0$ and $t\geq 0$ has to be considered. The considered post-quench and pre-quench dimensionless interaction parameters are $U_{\mathrm{f}}n(0)/J=0.5$ and $U_{\mathrm{i}}n(0)/J=2$ respectively. The dimensionless dissipation strength is given by $\gamma/J=0.1$.
@@
Figure 5: Quench spectral function $S_{k}(\omega)$ associated to (top panel) $G_{2}(R,t)=\langle\hat{n}_{R}(t)\hat{n}_{0}(t)\rangle_{\mathrm{c}}$ the equal-time connected density-density correlation function (bottom panel) $G_{1}(R,t)=\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{t}-\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{0}$ the equal-time connected one-body correlation function. A weak sudden double quench is performed on the dissipation strength from $\gamma=0$ to $\gamma>0$ as well as on the two-body repulsive interaction strength $U$ from $U=U_{\mathrm{i}}$ to $U=U_{\mathrm{f}}$. The dashed-dotted magenta lines represent twice and minus twice the post-quench quasiparticle dispersion relation associated to the isolated BH chain confined in the SF-mean-field regime, i.e. $2\mathcal{E}_{k,\mathrm{f}}$ and $-2\mathcal{E}_{k,\mathrm{f}}$ respectively, see Eq. (37). The considered post-quench and pre-quench dimensionless interaction parameters are $U_{\mathrm{f}}n(0)/J=0.5$ and $U_{\mathrm{i}}n(0)/J=2$ respectively. The dimensionless dissipation strength is given by $\gamma/J=1$.
@@
We turn to the case of relatively large post-quench interaction strength $U_{\mathrm{f}}$ while considering small dissipations, i.e. $\gamma\ll 1$. According to the mean-field condition given by $\bar{n}\gg U/J$, for large enough $U_{\mathrm{f}}$, the 1D BH model at equilibrium is no longer confined in the mean-field regime of the SF phase but rather in the strongly-correlated regime. In what follows, we consider the specific case where $J=U_{\mathrm{f}}=\bar{n}=N(0)/L$. However, from Ref. despres2019 , it has been shown that the quasiparticle dispersion relation $\mathcal{E}_{k}$ for the BH chain confined in the SF-mean-field regime defined at Eq. (37) remains accurate far beyond its validity domain. Hence, even at relatively large post-quench interaction strength $U_{\mathrm{f}}$, we should be able to rely on the QS approach in order to get an accurate description of the low-lying excitation spectrum of the 1D BH model confined in the SF-strongly-correlated regime. On Fig. 6, the QSF associated to the density fluctuations characterized via the ETCCF $G_{2}$ and to the phase fluctuations via $G_{1}$ are compared to the theoretical expression of the quasiparticle dispersion relation $\mathcal{E}_{k,\mathrm{f}}$ initially valid in the SF-mean-field regime. As shown on both panels, we find a very good agreement between the low-lying excitation spectra found from the QSFs and $\mathcal{E}_{k,\mathrm{f}}$.
@@
Figure 6: Quench spectral function $S_{k}(\omega)$ associated to (top panel) $G_{2}(R,t)=\langle\hat{n}_{R}(t)\hat{n}_{0}(t)\rangle_{\mathrm{c}}$ the equal-time connected density-density correlation function (bottom panel) $G_{1}(R,t)=\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{t}-\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{0}$ the equal-time connected one-body correlation function. A weak sudden double quench is performed on the dissipation strength from $\gamma=0$ to $\gamma>0$ as well as on the two-body repulsive interaction strength $U$ from $U=U_{\mathrm{i}}$ to a relatively large post-quench interaction strength $U=U_{\mathrm{f}}$. The dashed-dotted magenta lines represent twice and minus twice the post-quench quasiparticle dispersion relation associated to the isolated BH chain confined in the SF-mean-field regime, i.e. $2\mathcal{E}_{k,\mathrm{f}}$ and $-2\mathcal{E}_{k,\mathrm{f}}$ respectively, see Eq. (37). The considered post-quench and pre-quench dimensionless interaction parameters are $U_{\mathrm{f}}n(0)/J=1$ and $U_{\mathrm{i}}n(0)/J=2$ respectively. The dimensionless dissipation strength is given by $\gamma/J=0.1$.
@@
III Experimental feasibility of the quench spectroscopy method
@@
[…]
The latter QS method to probe low-lying excitation spectra can be extended to dissipative many-body quantum fermionic and spin lattice chains, to a higher dimensionality of the lattice and to long-range interactions. Most importantly, this technique can be applied to experimental measurements obtained using quantum simulators based on ultra-cold atoms to engineer bosonic lattice models, neutral Rydberg atoms or trapped ions for the simulation of fermionic and spin lattice models. To conclude, QS can be seen as an alternative approach to the standard experimental techniques to measure excitation spectra such as the angle-resolved photoemission spectroscopy (ARPES)  stewart2008 or Bragg spectroscopy rey2005 .
IV Applicability of the quench spectroscopy to a non-Hermitian quantum lattice model